How do you calculate GDP with the expenditure approach if not all values are provided? Please
provide an example of how to answer a question which asks you to find GDP through the expenditure approach if all the values that are needed to answer the question are not provided.

Please do not answer with the definition of GDP and the expenditure approach. Do not answer this question if you do not know the answer.
Answer #1

I will illustrate a numerical example to make you understand the Expenditure method of measuring GDP( gross domestic product at market price)

Illustration------

$( trillion

Govt final Consumption expenditure 40

Private final Consumption Expenditure. 35

Net domestic fixed capital formation. 10

Consumption of fixed capital. 1

Closing stock. 17

Opening stock. 9

Exports. 7

Imports. 2

Net Factor income from abroad. 1

Net indirect taxes. 3

Lets solve this------------

As we know--------

Y= C+I + nx. ( We assume that national income of a country is equal to totsl Expenditure on goods and services)

WhereY= national income or GDP

* C Consumption expenditure or govt final Consumption expenditure ( GFCF) + private final Consumption expenditure (PFCF)

*I= investment Expenditure or GDCF or ( Gross domestic capital formation ) or { net domestic capital investment Expenditure( NDFCF) + Consumption of fixed capital ( Depreciation or D)+ change in stock}

Where change in stock= Opening stock - closing stock or ∆ S

To make simple ,I= NDFCF+ D + ∆S

*Nx = Exports- Imports = net exports

So, additing these three stars *,we get GDP

Now lets solve it-------

$( trillion)

PFCF. . 40

GFCF. 35

*C= ( Pfcf+ Gfcf). 75

NDFCF. 10

D( Consumption of fixed c). 1

∆S. ( 17-9). 8

* I. ( NDFCF+ D+ ∆S). 19

Exports. (X). 7

Imports. ( M). 2

* NX ( X- M ). 5

GDP. = C+I+ NX

75+19+5= 99

The information regarding net indirect taxes and net Factor income from abroad is used if we are to calculate GNP and GDP at factor cost.

So,GDP = 99 trillion Dollars

( hope it will help you)